# The lower $p$-series of analytic pro-$p$ groups and Hausdorff dimension

### v1

@@ Abstract.
Furthermore, we show that, if $G$ is a countably based pro-$p$ group with an open subgroup mapping onto the free abelian pro-$p$ group $\mathbb{Z}_{p}\oplus\mathbb{Z}_{p}$, then for every prescribed finite set $\{0,1\}\subseteq X\subseteq[0,1]$ there is a filtration series $\mathcal{S}$ such that $\operatorname{hspec}^{\mathcal{S}}(G)=X$; in particular, $\lvert\operatorname{hspec}^{\mathcal{S}}(G)\rvert$ is unbounded, as $\mathcal{S}$ runs through all filtration series of $G$ with $\lvert\operatorname{hspec}^{\mathcal{S}}(G)\rvert<\infty$.
Finally, we establish that finitely generated non-abelian direct products $G$ of free pro-$p$ groups have full Hausdorff spectrum $\operatorname{hspec}^{\mathcal{L}}(G)=[0,1]$ with respect to the lower $p$-series $\mathcal{L}$.
@@
Lower $p$-series, $p$-adic analytic groups, Hausdorff dimension, Hausdorff spectrum, free pro-$p$ groups
@@ 1. Introduction
Throughout $p$ denotes a prime, and $\mathbb{Z}_{p}$ is the ring of $p$-adic integers. We are interested in the lower $p$-series of $p$-adic analytic pro-$p$ groups and questions about Hausdorff dimension with respect to the lower $p$-series. To put our results into context, we also consider specific situations involving more general pro-$p$ groups and other filtration series.
@@
where we adopt the notation from [7, Sec. 1.2]. It is the fastest descending central series of $G$ with elementary abelian sections and it consists of open normal subgroups forming a base of neighbourhoods for the element $1$. The lower $p$-series plays a significant role in the study of pro-$p$ groups. For instance, $\bigoplus_{i=1}^{\infty}P_{i}(G)/P_{i+1}(G)$ naturally carries the structure of an $\mathbb{F}_{p}$-Lie algebra and can frequently be used to analyse the structure of $G$ or its finite quotients; compare with [16, Chap. VIII], [24, Chap. II] and [23, Sec. 5].
@@
is the kernel of the natural epimorphism from $\mathbb{Z}_{p}G$ onto the finite field $\mathbb{F}_{p}$ that sends each group element to $1$; it is convenient to refer to $\mathfrak{a}_{G}$, somewhat informally, as the $p$-augmentation ideal of the group ring $\mathbb{Z}_{p}G$. The powers $\mathfrak{a}_{G}^{\,i}$, $i\in\mathbb{N}_{0}$, of the $p$-augmentation ideal intersect in $\{0\}$ and induce a topology on $\mathbb{Z}_{p}G$ which gives rise to the completed group ring $\mathbb{Z}_{p}[\![G]\!]$; see [7, Sec. 7.1]. The descending series of open $\mathbb{Z}_{p}G$-submodules
@@
The above result yields an analogous description of the lower $p$-series of a $p$-adic analytic pro-$p$ group. We recall that, for every $p$-adic analytic pro-$p$ group $G$, there exists $j\in\mathbb{N}$ such that for every $i\in\mathbb{N}$ with $i\geq j$ the group $P_{i}(G)$ is uniformly powerful; see [7, Prop. 3.9 and Thm. 4.2]. Moreover, every minimal generating system $x_{1},\ldots,x_{d}$ of a uniformly powerful pro-$p$ group $U$ yields a product decomposition of $U=\overline{\langle x_{1}\rangle}\cdots\overline{\langle x_{d}\rangle}$ into $d$ procyclic subgroups, which in turn gives a system of $p$-adic coordinates $U\to\mathbb{Z}_{p}^{\,d}$, $x_{1}^{\,a_{1}}\cdots x_{d}^{\,a_{d}}\mapsto(a_{1},\ldots,a_{d})$ of ‘the second kind’; see [7, Thm. 4.9]. The dimension of $G$, denoted by $\dim(G)$, coincides with the rank of any uniformly powerful open subgroup of $G$.
@@
Even for well-behaved groups, such as $p$-adic analytic pro-$p$ groups $G$, not only the Hausdorff dimension of individual subgroups, but also qualitative features of the Hausdorff spectrum of $G$ are known to be sensitive to the choice of $\mathcal{S}$; see [22, Thm. 1.3]. In our setting, where $G$ is a finitely generated pro-$p$ group, the group-theoretically relevant choices of $\mathcal{S}$ are primarily: the $p$-power series $\mathcal{P}$, the iterated $p$-power series $\mathcal{P}^{*}$, the Frattini series $\mathcal{F}$, the dimension subgroup (or Zassenhaus) series $\mathcal{D}$ and, last but not least, the lower $p$-series $\mathcal{L}\colon P_{i}(G),\,i\in\mathbb{N}$. It was shown in [22, Prop. 1.5] that, if $G$ is a $p$-adic analytic pro-$p$ group, then the Hausdorff dimensions of $H\leq_{\mathrm{c}}G$ with respect to the series $\mathcal{P},\mathcal{F},\mathcal{D}$ coincide, and [5, Thm. 1.1] yields, for $\mathcal{S}$ any one of these series $\mathcal{P},\mathcal{F},\mathcal{D}$, that the spectrum
@@
It was seen in [22] that, contrary to indications in [5], the Hausdorff spectra of $p$-adic analytic pro-$p$ groups $G$ with respect to the lower $p$-series $\mathcal{L}$ follow a different pattern. Concretely, [22, Example 4.1] provides a family of $p$-adic analytic pro-$p$ groups $G$ such that $\lvert\operatorname{hspec}^{\mathcal{L}}(G)\rvert/\dim(G)$ is unbounded as $\dim(G)\to\infty$. It remained open whether $\operatorname{hspec}^{\mathcal{L}}(G)$ could even be infinite for some $p$-adic analytic groups $G$. Theorem 1.1 allows us to establish that $\operatorname{hspec}^{\mathcal{L}}(G)$ is finite for every $p$-adic analytic pro-$p$ group $G$ and to put in place further restrictions.
@@
Theorem 1.5 triggers more refined questions about Hausdorff spectra with respect to the lower $p$-series. Under additional assumptions on $G$, we obtain the following stronger results, which partly extend even to other filtration series.
Theorem 1.6.
Let $G$ be an infinite $p$-adic analytic pro-$p$ group.
If $G$ is just infinite, then every closed subgroup $H\leq_{\mathrm{c}}G$ has strong Hausdorff dimension $\operatorname{hdim}_{G}^{\mathcal{S}}(H)=\dim(H)/\dim(G)$ for every filtration series $\mathcal{S}$.
If $G$ is nilpotent, then every closed subgroup $H\leq_{\mathrm{c}}G$ has strong Hausdorff dimension $\operatorname{hdim}_{G}^{\mathcal{S}}(H)=\dim(H)/\dim(G)$ for $\mathcal{S}=\mathcal{L}$.
In the cases discussed above, the Hausdorff spectrum of $G$ with respect to the relevant filtration series is equal to $\Big{\{}\nicefrac{{\displaystyle\dim(H)}}{{\displaystyle\dim(G)}}\mid H\leq_{\mathrm{c}}G\Big{\}}$.
It remains an outstanding problem in the subject to understand, more generally, what renders the Hausdorff spectrum $\operatorname{hspec}^{\mathcal{S}}(G)$ of a finitely generated pro-$p$ group $G$ finite. In particular, it is open whether finiteness of the spectrum with respect to one of the standard filtration series $\mathcal{P},\mathcal{P}^{*},\mathcal{F},\mathcal{D},\mathcal{L}$ implies that $G$ is $p$-adic analytic; compare with [22, Prob. 1.4]. It is known that for every $p$-adic analytic pro-$p$ group $G$ whose abelianisation has torsion-free rank at least $2$, there is a filtration $\mathcal{S}$ such that $\operatorname{hspec}^{\mathcal{S}}(G)$ is infinite; see [22, Thm. 1.3]. This leads to the question [22, Prob. 1.6]: does there exist, for every $p$-adic analytic pro-$p$ group $G$, a uniform bound $b(G)$ for $\lvert\operatorname{hspec}^{\mathcal{S}}(G)\rvert$, as $\mathcal{S}$ runs through all filtration series of $G$ with $\lvert\operatorname{hspec}^{\mathcal{S}}(G)\rvert<\infty$. We show that the answer is typically negative, in the strongest possible way.
Theorem 1.7.
Let $G$ be a countably based pro-$p$ group that has an open subgroup mapping surjectively onto $\mathbb{Z}_{p}\oplus\mathbb{Z}_{p}$. Then for every finite subset $X\subseteq[0,1]$ with $\{0,1\}\subseteq X$ there exists a filtration series $\mathcal{S}$ such that $\operatorname{hspec}^{\mathcal{S}}(G)=X$.
Irrespective of whether or not we expect the finiteness of the Hausdorff spectrum to characterise $p$-adic analytic pro-$p$ groups, it is of interest to produce and study potential counterexamples to such a proposition. For the filtration series $\mathcal{F}$ and $\mathcal{D}$, it was shown in [9] that free pro-$p$ groups $F$ can be cancelled from the candidate list; in fact, non-abelian finite direct products $G$ of free pro-$p$ groups have full Hausdorff spectrum $\operatorname{hspec}^{\mathcal{F}}(G)=\operatorname{hspec}^{\mathcal{D}}(G)=[0,1]$. Using Lie-theoretic tools, we produce an analogous result for the lower $p$-series $\mathcal{L}$.
Theorem 1.8.
Let $G=F_{1}\times\ldots\times F_{r}$ be a finite direct product of finitely generated free pro-$p$ groups $F_{j}$, at least one of which is non-abelian. Then $G$ has full Hausdorff spectrum $\operatorname{hspec}^{\mathcal{L}}(G)=[0,1]$ with respect to the lower $p$-series $\mathcal{L}$.
It remains a challenge to compute the Hausdorff spectra with respect to the lower $p$-series for pro-$p$ groups of positive rank gradient, such as non-soluble Demushkin pro-$p$ groups. This would further complement the results in [9].
@@
Organisation. In Section 2, we study stratifications of the lower $p$-series and establish Theorems 1.1 and 1.2, as well as their corollaries. In Section 3 we turn to Hausdorff spectra with respect to the lower $p$-series: using the results from the previous section, we prove Theorems 1.5 and 1.6, for $p$-adic analytic pro-$p$ groups. In Section 4 we establish Theorem 1.7, yielding arbitrarily large finite Hausdorff spectra, and Theorem 1.8 about full Hausdorff spectra for finite direct products of free pro-$p$ groups.
@@ 2. Stratifications of the lower $p$-series
Let $G$ be a profinite group, and let $L$ be a $\mathbb{Z}_{p}$-lattice equipped with a right $G$-action. We are interested in situations, where the lower $p$-series $\lambda_{i}(L)$, $i\in\mathbb{N}_{0}$, of $L$, as defined in (1.1), satisfies $\bigcap_{i\in\mathbb{N}_{0}}\lambda_{i}(L)=\{0\}$ and hence gives a filtration series. This is the case if and only if $G$ acts unipotently on every principal congruence quotient $L/p^{j}L$, $j\in\mathbb{N}$, which in turn is true if and only if the action of $G$ on the $\mathbb{Z}_{p}$-lattice $L\cong\mathbb{Z}_{p}^{\,d}$ factors through a pro-$p$ subgroup of $\mathsf{GL}_{d}(\mathbb{Z}_{p})$. Since every closed subgroup of $\mathsf{GL}_{d}(\mathbb{Z}_{p})$ is topologically finitely generated, we may restrict without any true loss of generality to the situation that $G$ itself is a finitely generated pro-$p$ group. A theorem of Serre implies that every abstract action of the group $G$ on the group $L$ is continuous; see [7, Sec. 1]. Thus we arrive at the standard set-up which we use henceforth: $L$ is a $\mathbb{Z}_{p}$-lattice equipped with a right $G$-action, for a finitely generated pro-$p$ group $G$.
@@
Let $(x_{1},\ldots,x_{d})$ be the frame for a stratification of $\mathcal{S}$, with growth rate $(\xi_{1},\ldots,\xi_{d})$, say. Then there is a constant $c\in\mathbb{N}$ such that $\mathcal{S}$ is $c$-equivalent to the filtration series $\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}$, $i\in\mathbb{N}_{0}$. Recall that $\xi_{1}\leq\ldots\leq\xi_{d}$ and choose $j\in\mathbb{N}$ such that $\lfloor j\xi_{e}\rfloor+c<\lfloor j\xi_{e+1}\rfloor-c$ for all $e\in\{1,\ldots,d-1\}$ with $\xi_{e}<\xi_{e+1}$. This implies that for all $i\in\mathbb{N}$ with $i\geq j$,
@@
The following auxiliary result is a variation of [19, Prop. 4.3] and can be proved in the same way, using compactness.
@@
Denote by $L_{i}=\lambda_{i}(L)$, $i\in\mathbb{N}_{0}$, the terms of the lower $p$-series of $L$, and for $i\in\mathbb{N}_{0}$ put $u_{i}=u_{L}(L_{i})$. By Proposition 2.2 the $\mathbb{Z}_{p}G$-module $L$ has finite rigidity $r=r(L)$. Consequently, the lower $p$-series of $L$ is equivalent to the filtration series $\mathcal{S}\colon L\geq p^{u_{1}}L\geq p^{u_{2}}L\geq\ldots$, and it suffices to show that $\mathcal{S}$ is equivalent to a series of the form $p^{\lfloor i\xi\rfloor}L$, $i\in\mathbb{N}_{0}$, for suitable $\xi\in[\nicefrac{{1}}{{d}},1]\cap\mathbb{Q}$.
@@
defined by the dictum $p^{u_{i}}x_{k}\mapsto z_{k}$ for $1\leq k\leq d$. For each $i\in\mathbb{N}$, we have $\ell_{L}(L_{i})\leq u_{i}+r$, hence $p^{u_{i}+r}L\leq L_{i}\leq p^{u_{i}}L$, and we consider $R_{i}=L_{i}\varphi_{i}$.
@@
$c=1+\max\big{\{}\lvert u_{i}-i\xi\rvert\bigm{|}0\leq i\leq j+m-1\big{\}}.$
@@
and the filtration series $\mathcal{S}$ is $c$-equivalent to $\mathcal{S}^{*}\colon p^{\lfloor i\xi\rfloor}L$, $i\in\mathbb{N}_{0}$. Finally, we notice that $L=L_{0}\supsetneqq L_{1}\supsetneqq\ldots$ yields the trivial lower bound $\lvert L:L_{i}\rvert\geq p^{i}$ and hence
@@
The $\mathbb{Z}_{p}$-lattice $L$ decomposes as the direct sum of its sublattices $M$ and $N=\bigoplus_{k=e+1}^{d}\mathbb{Z}_{p}x_{k}$. Choose $c_{1}\in\mathbb{N}$ such that the lower $p$-series $L_{i}=\lambda_{i}(L)$, $i\in\mathbb{N}_{0}$, is $c_{1}$-equivalent to the series $\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}$, $i\in\mathbb{N}_{0}$. Then
$L_{i}\cap N\subseteq p^{\lfloor i\xi_{e+1}\rfloor-c_{1}}N\qquad\text{for all $i\in\mathbb{N}_{0}$,}$ (2.1)
@@
For a contradiction, assume that $M$ is not $G$-invariant, i.e. not a $\mathbb{Z}_{p}G$-submodule. Then there are elements $\alpha_{1},\ldots,\alpha_{e}\in\mathbb{Z}_{p}G$ and $c_{2}\in\mathbb{N}$ such that $y=\sum_{k=1}^{e}x_{k}\alpha_{k}\in N\smallsetminus p^{c_{2}}N$. For every $i\in\mathbb{N}$, we have $p^{\lfloor i\xi_{e}\rfloor+c_{1}}M\subseteq L_{i}$ and hence $p^{\lfloor i\xi_{e}\rfloor+c_{1}}y\in(L_{i}\cap N)\smallsetminus p^{\lfloor i\xi_{e}\rfloor+c_{1}+c_{2}}N$. Thus (2.1) yields $\lfloor i\xi_{e}\rfloor+c_{1}+c_{2}>\lfloor i\xi_{e+1}\rfloor-c_{1}$. Since $\xi_{e}<\xi_{e+1}$ this yields the desired contradiction for sufficiently large $i\in\mathbb{N}$.
@@
It is clear that, for all $i\in\mathbb{N}_{0}$,
$M_{i}\subseteq L_{i}\cap M\subseteq\left(\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor-c_{1}}\mathbb{Z}_{p}\,x_{k}\right)\cap M\subseteq p^{-b}\,\bigoplus_{k=1}^{e}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}\,x_{k},$
and it suffices to show that, for all $i\in\mathbb{N}_{0}$ with $i\geq b_{1}$,
@@
Suppose that we can arrange for $M$ to be $G$-invariant. Then $L$ is a direct sum $M\oplus N$ of $\mathbb{Z}_{p}G$-modules and its lower $p$-series decomposes accordingly: $\lambda_{i}(L)=\lambda_{i}(M)\oplus\lambda_{i}(N)$ for $i\in\mathbb{N}_{0}$. If both $\dim_{\mathbb{Z}_{p}}M$ and $\dim_{\mathbb{Z}_{p}}N$ are less than $d$, our claim follows by induction, while in the remaining case, $L=N$ and $M=\{0\}$, the proof can be finished by appealing to Proposition 2.3.
Now suppose that neither $M$ nor any other $\mathbb{Z}_{p}$-sublattice $M^{\prime}$ of $L$, which complements $N$ to yield an open sublattice $M^{\prime}\oplus N$ of $L$, is $G$-invariant. Let us fix some additional notation. We write
@@
In fact, we arrange that among all possible choices – for the complement $M$ for $N$ in $L$ (initially up to finite index), for the stratification based on the frame $(\mathbf{x}_{1},\ldots,\mathbf{x}_{e})$ and for the representatives $x_{1},\ldots,x_{e}$ in $L$ – we make one that renders $\tilde{e}$ as large as possible. In addition we choose $\hat{e}\in\{\tilde{e}+1,\ldots,e\}$ maximal with $\xi_{\hat{e}}=\xi_{\tilde{e}+1}$, and for short we put
@@
Recall that $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}N$ is a simple $\mathbb{Q}_{p}G$-module. Proposition 2.2 shows that the non-zero $G$-invariant $\mathbb{Z}_{p}$-lattice $\big{(}\tilde{M}+\tilde{T}_{\mathbf{v}}.\mathbb{Z}_{p}G\big{)}\cap N$ is open in $N$ of lower level $l_{\mathbf{v}}\in\mathbb{N}_{0}$, say.

### v2

@@ Abstract.
Lower $p$-series, $p$-adic analytic groups, Hausdorff dimension, Hausdorff spectrum
@@ 1. Introduction
Throughout $p$ denotes a prime, and $\mathbb{Z}_{p}$ is the ring of $p$-adic integers. We are interested in the lower $p$-series of $p$-adic analytic pro-$p$ groups and questions about Hausdorff dimension with respect to the lower $p$-series.
@@
where we adopt the notation from [5, Sec. 1.2]. It is the fastest descending central series of $G$ with elementary abelian sections and it consists of open normal subgroups forming a base of neighbourhoods for the element $1$. The lower $p$-series plays a significant role in the study of pro-$p$ groups. For instance, $\bigoplus_{i=1}^{\infty}P_{i}(G)/P_{i+1}(G)$ naturally carries the structure of an $\mathbb{F}_{p}$-Lie algebra and can frequently be used to analyse the structure of $G$ or its finite quotients; compare with [13, Chap. VIII], [19, Chap. II] and [18, Sec. 5].
@@
is the kernel of the natural epimorphism from $\mathbb{Z}_{p}G$ onto the finite field $\mathbb{F}_{p}$ that sends each group element to $1$; it is convenient to refer to $\mathfrak{a}_{G}$, somewhat informally, as the $p$-augmentation ideal of the group ring $\mathbb{Z}_{p}G$. The powers $\mathfrak{a}_{G}^{\,i}$, $i\in\mathbb{N}_{0}$, of the $p$-augmentation ideal intersect in $\{0\}$ and induce a topology on $\mathbb{Z}_{p}G$ which gives rise to the completed group ring $\mathbb{Z}_{p}[\![G]\!]$; see [5, Sec. 7.1]. The descending series of open $\mathbb{Z}_{p}G$-submodules
@@
The above result yields an analogous description of the lower $p$-series of a $p$-adic analytic pro-$p$ group. We recall that, for every $p$-adic analytic pro-$p$ group $G$, there exists $j\in\mathbb{N}$ such that for every $i\in\mathbb{N}$ with $i\geq j$ the group $P_{i}(G)$ is uniformly powerful; see [5, Prop. 3.9 and Thm. 4.2]. Moreover, every minimal generating system $x_{1},\ldots,x_{d}$ of a uniformly powerful pro-$p$ group $U$ yields a product decomposition of $U=\overline{\langle x_{1}\rangle}\cdots\overline{\langle x_{d}\rangle}$ into $d$ procyclic subgroups, which in turn gives a system of $p$-adic coordinates $U\to\mathbb{Z}_{p}^{\,d}$, $x_{1}^{\,a_{1}}\cdots x_{d}^{\,a_{d}}\mapsto(a_{1},\ldots,a_{d})$ of ‘the second kind’; see [5, Thm. 4.9]. The dimension of $G$, denoted by $\dim(G)$, coincides with the rank of any uniformly powerful open subgroup of $G$.
@@
Even for well-behaved groups, such as $p$-adic analytic pro-$p$ groups $G$, not only the Hausdorff dimension of individual subgroups, but also qualitative features of the Hausdorff spectrum of $G$ are known to be sensitive to the choice of $\mathcal{S}$; see [17, Thm. 1.3]. In our setting, where $G$ is a finitely generated pro-$p$ group, the group-theoretically relevant choices of $\mathcal{S}$ are primarily: the $p$-power series $\mathcal{P}$, the iterated $p$-power series $\mathcal{P}^{*}$, the Frattini series $\mathcal{F}$, the dimension subgroup (or Zassenhaus) series $\mathcal{D}$ and, last but not least, the lower $p$-series $\mathcal{L}\colon P_{i}(G),\,i\in\mathbb{N}$. It was shown in [17, Prop. 1.5] that, if $G$ is a $p$-adic analytic pro-$p$ group, then the Hausdorff dimensions of $H\leq_{\mathrm{c}}G$ with respect to the series $\mathcal{P},\mathcal{F},\mathcal{D}$ coincide. Furthermore, [3, Thm. 1.1] yields, for $\mathcal{S}$ any one of these series $\mathcal{P},\mathcal{F},\mathcal{D}$, that the spectrum
@@
It was seen in [17] that, contrary to indications in [3], the Hausdorff spectra of $p$-adic analytic pro-$p$ groups $G$ with respect to the lower $p$-series $\mathcal{L}$ follow a different pattern. Concretely, [17, Example 4.1] provides a family of $p$-adic analytic pro-$p$ groups $G$ such that $\lvert\operatorname{hspec}^{\mathcal{L}}(G)\rvert/\dim(G)$ is unbounded as $\dim(G)\to\infty$. It remained open whether $\operatorname{hspec}^{\mathcal{L}}(G)$ could even be infinite for some $p$-adic analytic groups $G$. Theorem 1.1 allows us to establish that $\operatorname{hspec}^{\mathcal{L}}(G)$ is in fact finite for every $p$-adic analytic pro-$p$ group $G$ and to put in place further restrictions.
@@
Theorem 1.5 triggers more refined questions about Hausdorff spectra with respect to the lower $p$-series. We address some of these in a separate paper. For instance, one outstanding problem is to understand what renders the Hausdorff spectrum of a finitely generated pro-$p$ group finite. Using Lie-theoretic tools we show that non-abelian finite direct products of free pro-$p$ groups have full Hausdorff spectrum $[0,1]$ with respect to the lower $p$-series, in analogy to [7].
@@
Organisation. In Section 2, we study stratifications of the lower $p$-series and establish Theorems 1.1 and 1.2, as well as their corollaries. In Section 3 we turn to Hausdorff spectra with respect to the lower $p$-series: using the results from the previous section, we prove Theorem 1.5.
@@ 2. Stratifications of the lower $p$-series
Let $G$ be a profinite group, and let $L$ be a $\mathbb{Z}_{p}$-lattice equipped with a right $G$-action. We are interested in situations, where the lower $p$-series $\lambda_{i}(L)$, $i\in\mathbb{N}_{0}$, of $L$, as defined in (1.1), satisfies $\bigcap_{i\in\mathbb{N}_{0}}\lambda_{i}(L)=\{0\}$ and hence gives a filtration series. This is the case if and only if $G$ acts unipotently on every principal congruence quotient $L/p^{j}L$, $j\in\mathbb{N}$, which in turn is true if and only if the action of $G$ on the $\mathbb{Z}_{p}$-lattice $L\cong\mathbb{Z}_{p}^{\,d}$ factors through a pro-$p$ subgroup of $\mathsf{GL}_{d}(\mathbb{Z}_{p})$. Since every closed subgroup of $\mathsf{GL}_{d}(\mathbb{Z}_{p})$ is topologically finitely generated, we may restrict without any true loss of generality to the situation that $G$ itself is a finitely generated pro-$p$ group. A theorem of Serre implies that every abstract action of the group $G$ on the group $L$ is continuous; see [5, Sec. 1]. Thus we arrive at the standard set-up which we use henceforth: $L$ is a $\mathbb{Z}_{p}$-lattice equipped with a right $G$-action, for a finitely generated pro-$p$ group $G$.
@@
Let $(x_{1},\ldots,x_{d})$ be the frame for a stratification of $\mathcal{S}$, with growth rate $(\xi_{1},\ldots,\xi_{d})$, say. Then there is a constant $c\in\mathbb{N}$ such that $\mathcal{S}$ is $c$-equivalent to the filtration series $\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}$, $i\in\mathbb{N}_{0}$. Recall that $\xi_{1}\leq\ldots\leq\xi_{d}$ and choose $j\in\mathbb{N}$ such that $\lfloor j\xi_{e}\rfloor+c<\lfloor j\xi_{e+1}\rfloor-c$ for all $e\in\{1,\ldots,d-1\}$ with $\xi_{e}<\xi_{e+1}$. By definition of stratification, this implies that for all $i\in\mathbb{N}$ with $i\geq j$,
@@
For the trivial lattice $L=\{0\}$ the formalism yields $r(L)=-\infty$.
The following auxiliary result is a variation of [14, Prop. 4.3] and can be proved in the same way, using compactness.
@@
Denote by $L_{i}=\lambda_{i}(L)$, $i\in\mathbb{N}_{0}$, the terms of the lower $p$-series of $L$, and for $i\in\mathbb{N}_{0}$ put $u_{i}=u_{L}(L_{i})$. By Proposition 2.2 the $\mathbb{Z}_{p}G$-module $L$ has finite rigidity $r=r(L)$. Consequently, the lower $p$-series of $L$ is $r$-equivalent to the filtration series $\mathcal{S}\colon L\geq p^{u_{1}}L\geq p^{u_{2}}L\geq\ldots$, and it suffices to show that $\mathcal{S}$ is equivalent to a series of the form $p^{\lfloor i\xi\rfloor}L$, $i\in\mathbb{N}_{0}$, for suitable $\xi\in[\nicefrac{{1}}{{d}},1]\cap\mathbb{Q}$.
@@
defined by the dictum $p^{u_{i}}x_{k}\mapsto z_{k}$ for $1\leq k\leq d$. For each $i\in\mathbb{N}$, we have $\ell_{L}(L_{i})\leq u_{i}+r$, hence $p^{u_{i}+r}L\leq L_{i}\leq p^{u_{i}}L$ by the definition of $u_{i}$ and $r$, and we consider $R_{i}=L_{i}\varphi_{i}$.
@@
$c=\bigl{\lceil}1+\max\big{\{}\lvert u_{i}-i\xi\rvert\bigm{|}0\leq i\leq j+m-1\big{\}}\bigr{\rceil}\in\mathbb{N}.$
@@
and the filtration series $\mathcal{S}$ is $c$-equivalent to $\mathcal{S}^{*}\colon p^{\lfloor i\xi\rfloor}L$, $i\in\mathbb{N}_{0}$. Furthermore, we notice that $L=L_{0}\supsetneqq L_{1}\supsetneqq\ldots$ yields the trivial lower bound $\lvert L:L_{i}\rvert\geq p^{i}$. Combining these two results, we arrive at
@@
The $\mathbb{Z}_{p}$-lattice $L$ decomposes as the direct sum of its sublattices $M$ and $N=\bigoplus_{k=e+1}^{d}\mathbb{Z}_{p}x_{k}$. Choose $c_{1}\in\mathbb{N}$ such that the lower $p$-series $L_{i}=\lambda_{i}(L)$, $i\in\mathbb{N}_{0}$, is $c_{1}$-equivalent to the series $\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}$, $i\in\mathbb{N}_{0}$. Then $\xi_{e+1}=\min\{\xi_{e+1},\ldots,\xi_{d}\}$ implies
$L_{i}\cap N\subseteq\left(\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor-c_{1}}\mathbb{Z}_{p}x_{k}\right)\cap N\subseteq p^{\lfloor i\xi_{e+1}\rfloor-c_{1}}N\quad\text{for all $i\in\mathbb{N}_{0}$,}$ (2.1)
@@
For a contradiction, assume that $M$ is not $G$-invariant, i.e. not a $\mathbb{Z}_{p}G$-submodule. Then there are elements $\alpha_{1},\ldots,\alpha_{e}\in\mathbb{Z}_{p}G$ and $c_{2}\in\mathbb{N}$ such that $y=\sum_{k=1}^{e}x_{k}\alpha_{k}\in N\smallsetminus p^{c_{2}}N$. For every $i\in\mathbb{N}$, we deduce from $\xi_{e}=\max\{\xi_{1},\ldots,\xi_{e}\}$ that $p^{\lfloor i\xi_{e}\rfloor+c_{1}}M\subseteq L_{i}$ and hence
$p^{\lfloor i\xi_{e}\rfloor+c_{1}}y\in(L_{i}\cap N)\smallsetminus p^{\lfloor i\xi_{e}\rfloor+c_{1}+c_{2}}N.$
Thus (2.1) yields $\lfloor i\xi_{e}\rfloor+c_{1}+c_{2}>\lfloor i\xi_{e+1}\rfloor-c_{1}$. Since $\xi_{e}<\xi_{e+1}$ this yields the desired contradiction for sufficiently large $i\in\mathbb{N}$.
@@
From the choice of $c_{1}$ and from $b\geq c_{1}$ we deduce that, for all $i\in\mathbb{N}_{0}$,
$M_{i}\subseteq L_{i}\cap M\subseteq\left(\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor-c_{1}}\mathbb{Z}_{p}\,x_{k}\right)\cap M\subseteq p^{-b}\,\bigoplus_{k=1}^{e}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}\,x_{k}.$
Thus it suffices to show that, for all $i\in\mathbb{N}_{0}$ with $i\geq b_{1}$,
@@
Suppose that we can arrange for $M$ to be $G$-invariant. Then $L$ is a direct sum $M\oplus N$ of $\mathbb{Z}_{p}G$-modules and its lower $p$-series decomposes accordingly: $\lambda_{i}(L)=\lambda_{i}(M)\oplus\lambda_{i}(N)$ for $i\in\mathbb{N}_{0}$. If both $\dim_{\mathbb{Z}_{p}}M$ and $\dim_{\mathbb{Z}_{p}}N$ are less than $d$, our claim follows by induction, while in the remaining case, $L=N$ and $M=\{0\}$, the proof can be finished by appealing to Proposition 2.3. Using Lemma 2.6, the same argument can be applied if $N$ admits a $G$-invariant almost complement $M^{\prime}$ in $L$, i.e. a $G$-invariant $\mathbb{Z}_{p}$-sublattice $M^{\prime}$ which complements $N$ to yield an open sublattice $M^{\prime}\oplus N$ of $L$.
Now suppose that $N$ has no almost complement in $L$. Let us fix some additional notation. We write
@@
In fact, we arrange that among all possible choices – for $M$ among all possible almost complements for $N$ in $L$ (with a subsequent reduction to $L=M\oplus N$), for the stratification based on the frame $(\mathbf{x}_{1},\ldots,\mathbf{x}_{e})$ and for the representatives $x_{1},\ldots,x_{e}$ in $L$ – we make one that renders $\tilde{e}$ as large as possible. In addition we choose $\hat{e}\in\{\tilde{e}+1,\ldots,e\}$ maximal with $\xi_{\hat{e}}=\xi_{\tilde{e}+1}$, and for short we put
@@
Recall that $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}N$ is a simple $\mathbb{Q}_{p}G$-module. Hence the non-zero $G$-invariant $\mathbb{Z}_{p}$-lattice $\big{(}\tilde{M}+\tilde{T}_{\mathbf{v}}.\mathbb{Z}_{p}G\big{)}\cap N$ is open in $N$ of lower level $l_{\mathbf{v}}\in\mathbb{N}_{0}$, say; compare with Proposition 2.2 and the paragraph above it.